OpenStudy (anonymous):

If a solid has 8 vertices and 16 edges, how many faces does it have

OpenStudy (rational):

you may use euler's polyhedron formula\[V-E+F=2\]

OpenStudy (anonymous):

So, 8-16+? would be like that

OpenStudy (rational):

Yes plugin the numbers and solve \(F\)

OpenStudy (anonymous):

A. 10 faces B. 12 faces C. 16 faces D. 20 faces

OpenStudy (rational):

\[8-16+F = 2\\-8 +F = 2\\F = 10\]

OpenStudy (anonymous):

thanks

OpenStudy (rational):

yw